Jason is 21 years older than his sister Sarah. In 5 years he will be twice as old as Sarah. How old is each of them now?
Answer by Maths68(1474)

Let
Present age of Jason = j
Present age of Sarah = s


Jason is 21 years older than his sister Sarah.
j=s+21.............(1)
In 5 years their age will be
j+5=s+5
then he will be twice as old as Sarah.
j+5=2(s+5)
j+5=2s+10
j=2s+10-5
j=2s+5..............(2)
Put the value of j from (1) to (2)
s+21=2s+5
s-2s=5-21
-s=-16
Multiply by -1 both sides
s=16
Put the value of s in (1)
j=s+21
j=16+21
j=37
Present age of Jason = j = 37 years old
Present age of Sarah = s = 16 years old